# Enhancing Visual Reasoning with Autonomous Imagination in Multimodal Large Language Models

### v1

@@ Abstract
There have been recent efforts to extend the Chain-of-Thought (CoT) paradigm to Multimodal Large Language Models (MLLMs) by finding visual clues in the input scene, advancing the visual reasoning ability of MLLMs. However, current approaches are specially designed for the tasks where clue finding plays a major role in the whole reasoning process, leading to the difficulty in handling complex visual scenes where clue finding does not actually simplify the whole reasoning task. To deal with this challenge, we propose a new visual reasoning paradigm enabling MLLMs to autonomously modify the input scene to new ones based on its reasoning status, such that CoT is reformulated as conducting simple closed-loop decision-making and reasoning steps under a sequence of imagined visual scenes, leading to natural and general CoT construction. To implement this paradigm, we introduce a novel plug-and-play imagination space, where MLLMs conduct visual modifications through operations like focus, ignore, and transform based on their native reasoning ability without specific training. We validate our approach through a benchmark spanning dense counting, simple jigsaw puzzle solving, and object placement, challenging the reasoning ability beyond clue finding. The results verify that while existing techniques fall short, our approach enables MLLMs to effectively reason step by step through autonomous imagination.
Project page: https://future-item.github.io/autoimagine-site.
Figure 1: Our autonomous imagination method empowers advanced MLLMs to engage in iterative imaginative reasoning, enabling them to address previously unsolvable tasks without additional training or fine-tuning.
@@ 1 Introduction
Recently, Multimodal Large Language Models (MLLMs) have achieved impressive advancements in the domain of visual understanding [38, 31, 3, 52, 75, 7, 36, 2]. Despite these strides, MLLMs still face challenges when tackling complex visual reasoning tasks, often producing erroneous outputs and hallucinations. Recent studies have promoted Chain-of-Thought (CoT) reasoning [55] in MLLMs based on the paradigm of visual clue finding, effectively advancing their reasoning abilities [22, 58, 5, 41, 45, 73]. The central idea lies in constructing CoT steps as letting MLLMs find visual clues by adding visual markers, such as bounding boxes, in the given input visual scene, highlighting the key regions to perceive and understand. In the tasks where clue finding significantly simplifies the whole reasoning task, these approaches achieve promising results.
However, there exist three fundamental obstacles for these methods in handling complex visual scenes, where designing CoT based on clue finding does not actually simplify the reasoning task. 1) Adding visual markers could also lead to an increase in visual noise in the input scene. Such visual noise could lead to more serious hallucinations instead of alleviating them. We show an example in Fig. 5(c) from the dense counting task to illustrate this phenomenon. 2) The process of clue finding does not necessarily lead to the decrease of hardness for single-step reasoning. For example, consider the task of jigsaw puzzle solving shown in Fig. 6, which is to identify the position of a target piece by considering its relationship to all other pieces. Note that identifying one key piece as a trustworthy visual clue also requires considering the complex surrounding relationships. In the worst case, this single step has a similar complexity to the whole task. 3) The reasoning task can involve a significant workload beyond clue finding. As shown in the object placement task in Fig. 4, the reasoning problem challenges the global understanding of object positions and relationships instead of identifying individual visual clues.
On the other hand, inspired by human reasoning, a general way for breaking complex visual reasoning into simpler steps is to gradually modify the input scene, and perform visual reasoning throughout this modification process. For example, for jigsaw puzzle solving, if the CoT process is defined as successively creating the virtual scenes of putting the target piece into one of the candidate positions and justifying its correctness, the whole task is successfully broken into simpler steps. Built on this insight, we introduce a novel visual reasoning paradigm in which MLLMs autonomously modify visual content based on the input scene to create a sequence of imagined scenes, and deal with one-step reasoning tasks based on them, all with its native reasoning ability. Given that existing MLLMs are not trained for such imaginative operations, implementing this paradigm presents a substantial challenge.
Our solution is to introduce plug-and-play imagination spaces, in which MLLMs autonomously choose to execute modular operations such as focus, ignore, and transform. These operations allow the model to systematically restructure the initially unorganized visual scene, generating a sequence of modified scenes that support multiple rounds of CoT reasoning. This iterative reconfiguration process transforms complex visual tasks into digestible sub-steps aligned with the model’s single-step reasoning capability. Through a closed-loop cycle of imagination and single-step reasoning, MLLMs progressively refine their understanding and derive accurate conclusions—all without requiring additional training or fine-tuning themselves.
To better challenge MLLMs’ visual reasoning ability, we introduce the Intuitive Visual Reasoning Benchmark (IVRB), which focuses on three typical tasks: dense counting, simple jigsaw puzzle solving, and object placement. These tasks represent challenges that are natural and intuitive for humans. More importantly, they all require reasoning beyond clue finding as discussed previously, outside of what current benchmarks address. Through both quantitative and qualitative experiments, we demonstrate that even though existing techniques fall short, our approach significantly advances the ability of MLLMs to handle these challenging tasks, expanding the boundaries of the visual reasoning ability of MLLMs through autonomous imagination.
@@ 2 Related Work
2.1 Multimodal Large Language Models
MLLMs have evolved in visual reasoning tasks, initially leveraging domain-specific expert models like HuggingGPT [47], MM-REACT [61], and VisualChatGPT [56]. The focus later shifted toward training LLMs with an adaptor for the other modal, such as LLaVA [31], BLIP-2 [24, 23], and MoVA [76]. Many recent MLLMs now exhibit native visual understanding through training on text-image pairs, with fine-tuning on question-answering tasks [31, 3, 52, 75, 7, 36, 2], including the state-of-the-art closed-source MLLM GPT-4o [38].
@@ 2.2 Visual Reasoning in MLLMs
Numerous studies explore reasoning paradigms in natural language utilizing their in-context learning ability [4], showing that Large Language Models (LLMs) improve through reasoning-based outputs [55, 21]. Subsequent works further enhance reasoning through in-context learning by improving the in-context sample selection [44, 35, 69, 10, 53, 27]. Recently, notable advancements in OpenAI’s o1 [39] have demonstrated that through scaling LLMs for inference-time reasoning before answering, LLMs can be greatly enhanced in their abilities.
With advancements in reasoning within text modalities, substantial efforts are now focused on enabling reasoning in visual tasks to enhance performance. Existing MLLMs often face limitations in direct visual perception and are prone to generating answers with hallucinations [70]. Initial approaches employed attention mechanisms to enhance question-answering capabilities [74], and subsequent solutions include strengthening reasoning in visual modality by training [29, 51] or employing auxiliary knowledge [37]. Additionally, researchers have proposed effective visual prompting methods to refine the focus of MLLMs [16, 28, 12, 68, 67, 37, 71, 65, 63]. See [57] for the recent comprehensive survey.
Attempts have been made to incorporate visual reasoning by generating visual promptings during the reasoning steps in CoT. Based on pioneer studies [70, 71, 66, 40], recent approaches build CoT by enabling MLLMs to conduct visual clue finding either through direct model training [22, 58, 5, 41, 45] or by utilizing plug-and-play visual processing models [73]. This paradigm strengthens reasoning by filtering extraneous information and emphasizing essential visual elements, allowing models to leverage visual prompts effectively without human intervention. However, they remain difficult to handle challenging reasoning tasks beyond clue finding as discussed in Sec. 1.
@@
In closed-world action planning where the visual scene is pre-processed and structured, various imagination techniques have been explored, such as using video generation models to simulate control processes as references [1, 9] or employing image generation to visualize target goals [72]. Methods that improve an LLM’s understanding of the current state, either through textual descriptions or visualizations, have shown to enhance decision-making in closed-world sequential action planning as well [32, 59]. However, these approaches are limited to closed-world settings, where imagination or visualization is primarily used to align the closed-world state space with MLLMs. Although these techniques have shown promising results in closed environments, they are designed specifically for constrained settings and cannot be directly extended to open-world contexts, where the visual scene is primitive and unstructured. In open worlds, the state space is undefined, and the solution space is vast. Leveraging generative models that interpret text prompts and operate in open-world scenarios could be a potential approach. However, such an advanced world model is yet to be developed.
@@
The closest studies to ours are from the field of robotics, where recent methods utilize the ability of MLLMs to perform object manipulation [18, 8]. These methods also create a virtual space and use MLLMs as evaluators to judge whether the object’s final state matches the instruction, in order to generate a final state where the object should be moved to. However, current approaches adopt random sampling to conduct exhaustive searches in the virtual environment, which is not feasible when the possible state space is large. We implement this method under visual reasoning tasks in the experiments to verify this argument.
2.3 Visual Reasoning Benchmarks for MLLMs
Following the rapid development of MLLMs, various benchmarks are proposed to evaluate their visual reasoning abilities from different aspects. See [25] for a recent comprehensive survey. In this section, we focus on discussing two categories of benchmarks, multi-step reasoning and object hallucination, which are closest to the major purpose of our work: extending the boundary of the CoT reasoning ability for state-of-the-art MLLMs.
Multi-step reasoning benchmarks. CoT reasoning involves breaking complicated tasks into multiple simpler steps. There have been several benchmarks proposed for evaluating such kind of multi-step reasoning ability. Some of them focus on tasks of math [34], logic [60], and video [54], where our current method is not applicable. On the other hand, the CoM benchmark [41] evaluates the ability to conduct chain of manipulations for fetching detailed evidence from inputs, and the Visual CoT benchmark [45] evaluates the visual clue finding ability. Both these two benchmarks focus on evaluating relatively small-scale MLLMs, requiring specific training to conduct CoT reasoning. Note that our target is to strengthen state-of-the-art large-scale MLLMs, such as GPT-4o, with their native reasoning ability, which is already strong enough to handle many general reasoning problems. Our IVRB benchmark is proposed for this purpose. We also note that our method may not be suitable for small-scale MLLMs, which do not have strong enough native reasoning ability to properly utilize the autonomous imagination operations.
Object hallucination benchmarks. Current MLLMs can suffer from hallucination by perceiving or generating non-existing objects in the input scene. A number of object hallucination benchmarks are proposed for analyzing this phenomenon [64, 30, 17, 50, 11, 48, 43, 49, 13, 33, 26, 6]. Even though such hallucination seems not to be directly related to CoT reasoning, we note that our paradigm can serve as a promising way to address this challenge, especially for the scenes where multiple objects exist. By the closed-loop iterative process of autonomous imagination, MLLMs can deal with the objects one by one as in our counting benchmark instead of handling them altogether, effectively reducing the possibility of hallucination. We conduct additional experiments on the recently proposed multi-object hallucination benchmark ROPE [6] to validate this argument. See Sec. A for details.
@@ 3 Method
3.1 Reasoning Paradigm
We propose a new paradigm, Autonomous Imagination, to realize CoT visual reasoning of MLLMs. We consider the following visual reasoning task where an unstructured input 2D or 3D visual scene is given, meaning that no semantic pre-processing, such as segmentation or grounding, is conducted initially. Denote the input scene as $o$. Given the input text prompt token $r_{0}$, the target of reasoning is to obtain the prediction probability $P(y|o,r_{0})$, where $y$ is the final output text token. We assume MLLMs take both images and text prompts as inputs. For 2D scenes, $o$ can be processed directly. For 3D scenes, we assume to utilize the images rendered from the proposed 3D imagination space, which will be introduced later in Sec. 3.2.
In general, for utilizing CoT, the reasoning problem of predicting $P(y|o,r_{0})$ can be transformed into intermediate steps of predicting $P(z_{t}|o,z_{t-1}),t\in 1,2,\dots,T$. Each $z_{t}=\{r_{t},c_{t}\}$ denotes the augmented textual prompts $r_{t}$ and visual information $c_{t}$ for deepening scene understanding and pushing forward reasoning. After the final reasoning step at $T$, the final output is set as $y=r_{T}$. The target transforms into predicting the following joint probability:
@@ 3.1 Reasoning Paradigm
where $z_{t_{1}:t_{2}}=\{z_{t_{1}},z_{t_{1}+1},\dots,z_{t_{2}}\}$ and $z_{0}=r_{0}$. What is essential in CoT design is to make each step simple enough to match the reasoning capability of MLLMs.
In our approach, we propose a new paradigm of CoT design. We introduce the imagination space, a virtual visual environment that can render imagined scenes $\hat{o}$, which are also images for MLLMs to perceive and elicit MLLMs to generate new augmented textual prompt $r$. Furthermore, we introduce a set of operators for modifying the imagined scenes into new ones. Denote by $v_{t}=\{\hat{o}_{t},r_{0:t}\}$ and $a_{t}$ the operator at step $t$. Our method transforms Eq. 1 into
@@
Eq. 3 shows that one-step reasoning is factorized into a decision function $\pi(a_{t}|v_{t-1})$, a scene modification function $\phi(\hat{o}_{t}|\hat{o}_{t-1},a_{t})$, and a reasoning function $\omega(r_{t}|\hat{o}_{t},r_{1:r-1})$. Given the current imagined scene $\hat{o}_{t-1}$ and augmented text prompt token $r_{t-1}$, the decision function chooses a specific scene modification operator $a_{t}$. The scene modification function then utilizes $a_{t}$ to update $\hat{o}_{t-1}$ into $\hat{o}_{t}$. The reasoning function finally updates $r_{t-1}$ into $r_{t}$ based on $\hat{o}_{t}$. The decision and reasoning functions rely purely on the native reasoning ability of MLLMs. The scene modification function is implemented inside the imagination space.
Comparing Eq. 1 and Eq. 2, our paradigm transforms CoT reasoning into a closed-loop decision-making and reasoning process. Furthermore, Eq. 3 indicates that the reasoning step at step $t$ is only dependent on $\hat{o}_{t-1},\hat{o}_{t}$ but independent of all previous scenes. In particular, the reasoning focuses on the current imagined scenes instead of the input scene $o$. We show that in many complex reasoning tasks, this paradigm effectively alleviates the visual noise issue and leads to effective single-step task decomposition.
Figure 2: An overview of our autonomous imagination method: The imagination space begins with an unstructured input scene and undergoes an iterative reasoning process. In each cycle, MLLMs first perceive the current state of the imagination space, select an operation to apply, and then reassess the updated imagination space. Upon completing this reasoning sequence, MLLMs generate an answer based on the cumulative context of the process and the final state of the imagination space.

### v3

# Autonomous Imagination: Closed-Loop Decomposition of Visual-to-Textual Conversion in Visual Reasoning for Multimodal Large Language Models

@@ Abstract
Under pure textual modality, Large Language Models (LLMs) have demonstrated remarkable success in complex reasoning tasks by decomposing them into simpler sub-problems. However, Multimodal Large Language Models (MLLMs) still struggle with some seemingly straightforward visual tasks, such as counting and solving jigsaw puzzles. We argue that these tasks challenge the ability of visual-to-textual conversion, where MLLMs convert visual information perceived from the input scene, to textual information for further reasoning and generating the answer. If the complexity of the visual input is beyond the perceptual capability of the MLLMs, without decomposing this conversion process, simply scaling inference-time reasoning cannot solve the task because it repeatedly encounters the same perceptual bottleneck. We propose an approach, autonomous imagination, to enable MLLMs to iteratively modify visual inputs (e.g. isolating objects, rearranging puzzle pieces) into intermediate visual states, decomposing visual-to-textual conversion into closed-loop visual modification steps. We show that, without any retraining, MLLMs can now solve tasks initially beyond their perceptual capability, highlighting that closed-loop visual modification can be an effective way of decomposing the visual reasoning task into solvable substeps. Project page: https://future-item.github.io/autoimagine-site/.
@@ 1 Introduction
Modern Large Language Models (LLMs) demonstrate exceptional reasoning capabilities through systematic task decomposition (Huang & Chang, 2023). By breaking complex problems into sequential subtasks, models achieve remarkable performance in textual reasoning domains (Yao et al., 2023; OpenAI, 2024b). More recently, Multimodal Large Language Models (MLLMs) (OpenAI, 2024a; Liu et al., 2023; Bai et al., 2023; Wang et al., 2023a) are gaining increasingly strong abilities in visual understanding. Inspired by work on visual prompting (Jiang et al., 2024; Lin et al., 2024; Hong et al., 2023; Zhang et al., 2023b; c; Mitra et al., 2023; Zheng et al., 2023; Zhang et al., 2024; Yu et al., 2024), recent methods enable MLLMs to perform visual reasoning by proactively adding visual prompts (such as bounding boxes) to the image, to reduce hallucination and handle challenging visual tasks without human intervention (Lai et al., 2023; Wu & Xie, 2023; Chen et al., 2023; Qi et al., 2024; Shao et al., 2024; Zhou et al., 2024b).
Despite these advancements, existing methods often struggle to make MLLMs natively handle intuitive visual reasoning tasks. When challenged with tasks such as jigsaw puzzle assembly (even when provided with explicit step-by-step instructions), or simply counting objects, state-of-the-art MLLMs consistently fail to reason through the entire process. A key observation is that, instead of challenging high-level reasoning abilities such as logical deduction, these tasks actually challenge the ability to perceive and understand complex visual scenes. Referring to the failure of existing MLLMs on them, we argue that ineffective task decomposition remains a fundamental cause. Reasoning in complex visual input requires the strong ability of visual-to-textual conversion, where MLLMs convert visual information perceived from the input scene to textual information for further reasoning and generating the answer. However, in current MLLM reasoning, the visual-to-textual conversion process remains an atomic operation rather than being divided into executable substeps. When the complexity of the visual scene is beyond the perceptual ability of MLLMs, without decomposing this conversion step, simply scaling inference-time reasoning cannot solve the task because it repeatedly encounters the same perceptual bottleneck.
We hypothesize that, similar to textual reasoning, where intermediate textual states are necessary, MLLMs need to generate intermediate images that gradually simplify the original scene or approach the target visual state. To test this, we propose enabling MLLMs to iteratively modify visual inputs (e.g., isolating objects, rearranging puzzle pieces) into intermediate visual states, decomposing visual-to-textual conversion into closed-loop visual modification steps, without any additional training of MLLMs. For simplicity, we refer to this approach as autonomous imagination. Technically, we implement the approach to support the conversion from an unstructured input visual scene into a structured representation and the visual operators that efficiently modify the visual content into new states (e.g., through transformation or erasing). This allows MLLMs to generate successive visual states that progressively simplify the original visual scene or gradually approach the target visual state in an autonomous fashion.
We conduct experiments under visual reasoning tasks that challenge the perception and understanding abilities of MLLMs under complex visual scenes, including counting, jigsaw puzzle solving, object placement, and multi-object hallucination. The results show that previously challenging visual reasoning tasks, which are beyond the perceptual capability of the MLLMs, can be effectively tackled natively by them equipped with our approach. This serves as empirical validation that a major limitation in visual reasoning stems from insufficient support for visual task decomposition, which we hope will inspire future research.
Figure 1: Our autonomous imagination approach empowers advanced MLLMs to engage in iterative imaginative reasoning, enabling them to address previously unsolvable tasks without additional training.
@@ 2 Related Work
2.1 Reasoning in LLMs
Numerous studies explore reasoning paradigms in natural language using their in-context learning ability (Brown et al., 2020), showing that Large Language Models (LLMs) improve through reasoning-based outputs (Wei et al., 2022; Kojima et al., 2022). Subsequent work further improves reasoning through in-context learning by improving in-context sample selection (Rubin et al., 2021; Lu et al., 2022; Zhang et al., 2022; Fu et al., 2022; Wang et al., 2022; Li et al., 2022b). Recently, notable advancements in OpenAI’s o1 (OpenAI, 2024b) have demonstrated that by scaling LLMs for inference-time reasoning before answering, LLMs can be greatly enhanced in reasoning. A key insight from these advances is that proper decomposition of the reasoning process into simpler subtasks is essential for successful reasoning (Wei et al., 2022).
@@ 2.2 Visual Reasoning in MLLMs
MLLMs have evolved in visual understanding ability, initially leveraging domain-specific expert models such as HuggingGPT (Shen et al., 2023), MM-REACT (Yang et al., 2023), and VisualChatGPT (Wu et al., 2023). The focus later shifted to training LLMs with an adapter for the other modal, such as LLaVA (Liu et al., 2023), BLIP-2 (Li et al., 2022a; 2023), and MoVA (Zong et al., 2024). Many recent MLLMs now exhibit native visual understanding through training in text-image pairs, with fine-tuning on question-answering tasks (Liu et al., 2023; Bai et al., 2023; Wang et al., 2023a; Zhu et al., 2023; Chen et al., 2022; Luo et al., 2023; Alayrac et al., 2022), including the state-of-the-art closed-source MLLM GPT-4o (OpenAI, 2024a).
In addition to reasoning within text modalities, substantial efforts are made in reasoning in visual tasks. Existing MLLMs often face limitations in direct visual perception and are prone to generating answers with hallucinations (Zhang et al., 2023d). Initial approaches used attention mechanisms to improve question-answering abilities (Zhou et al., 2021), and subsequent solutions include strengthening reasoning in the visual modality by training (Lin et al., 2023; Wang et al., 2023b) or employing auxiliary knowledge (Mitra et al., 2023). In addition, researchers have proposed effective visual prompting methods to refine the focus of MLLMs (Jiang et al., 2024; Lin et al., 2024; Hong et al., 2023; Zhang et al., 2023b; c; Mitra et al., 2023; Zheng et al., 2023; Zhang et al., 2024; Yu et al., 2024). See Wu et al. (2024a) for the recent comprehensive survey. Attempts have been made to utilize visual prompts in the Chain-of-Thought (CoT) paradigm (Wei et al., 2022). Based on pioneer studies (Zhang et al., 2023d; Zheng et al., 2023; Zhang et al., 2023a; Peng et al., 2023), recent approaches build CoT by enabling MLLMs to add visual prompts autonomously either through direct model training (Lai et al., 2023; Wu & Xie, 2023; Chen et al., 2023; Qi et al., 2024; Shao et al., 2024) or by utilizing external visual processing models (Zhou et al., 2024b). By using visual prompts (e.g., bounding boxes) to highlight key information in images, the visual-to-textual conversion step becomes easier, leading to a performance boost. However, even though visual prompts simplify the perception of input images, this paradigm still restricts to do visual-to-textual conversion only in one step, which is after all visual prompts are added in the input. As our experiments demonstrate, even when visual prompting is applied, visual-to-textual conversion remains a bottleneck in common visual reasoning tasks that are intuitive to humans, highlighting the necessity of effective task decomposition.
@@
We also discuss some recent work on action planning, which addresses a significantly different task from ours, while has also explored various imagination techniques, such as using video generation models to simulate control processes as references (Ajay et al., 2023; Du et al., 2023), employing image generation to visualize target goals (Zhou et al., 2024a), and improving an LLM’s understanding of the current state, either through textual descriptions or visualizations (Liu et al., 2022; Wu et al., 2024b). However, these approaches are limited to closed-world settings where the state and action spaces are predefined. Although these techniques have shown promising results in closed environments, they are designed specifically for constrained settings and cannot be directly extended to open-world contexts, where the visual scene is primitive and unstructured.
@@
The closest studies to ours are from the field of robotics, where recent methods utilize the ability of MLLMs to perform object manipulation (Kapelyukh et al., 2024; Ding et al., 2024). These methods also create a virtual space and use MLLMs as evaluators to judge whether the object’s final state matches the instruction, in order to generate a final state where the object should be moved to. However, current approaches adopt random sampling to conduct exhaustive searches in the virtual environment, which is not feasible when the possible state space is large. We implement this method in the experiments to verify this argument.
@@ 3 Method
3.1 Problem Formulation
We challenge MLLMs with visual reasoning tasks that are intuitive for humans: counting, jigsaw puzzle solving, object placement, and multi-object hallucination. For object placement, to capture a holistic top-down view, we reconstruct the input scene using 3D Gaussian Splatting (Kerbl et al., 2023) and render it from a top-down perspective. We directly feed the unstructured raw input scene into the model: no semantic pre-processing (e.g., segmentation or grounding) is applied initially, aligning with real-world settings where visual inputs are provided directly. We evaluate whether MLLMs can natively solve these tasks by proactively modifying the visual state on their own, with minimal reliance on external semantic visual understanding abilities (e.g., we avoid using object detection models for counting, which would constitute “cheating”).
3.2 Closed-Loop Reasoning Formulation
Figure 2: An overview of our autonomous imagination approach: The imagination space begins with an unstructured input scene and undergoes an iterative reasoning process. In each cycle, MLLMs first perceive the current state of the imagination space, select an operation to apply, and then reassess the updated imagination space. Upon completing this reasoning sequence, MLLMs generate an answer based on the cumulative context of the process and the final state of the imagination space.
To enable existing MLLMs to proactively modify the visual scene during reasoning, we propose a plug-and-play virtual space. In this framework, the MLLM is treated as a pre-trained policy that executes a sequence of operations. We refer to this space as the imagination space for simplicity. The MLLM can move a cursor to focus on an object, perform an operation on it (e.g., ignore or transform it), and then go back to use the cursor again until convergence. The imagination space continuously updates the MLLM with the latest visual state by dynamically re-rendering the scene.
Note that our only external model dependency is Segment Anything 2 (SAM2) (Ravi et al., 2024), which is invoked solely to isolate a visual element after the MLLM selects it via the cursor. Since our focus is on testing visual perception rather than high-level logical reasoning, we directly provide the MLLM with task-specific reasoning plans (e.g., “count by ignoring balls”) through system prompts. This mitigates occasional failures caused by the MLLM adopting bad reasoning strategies.
Specifically, we assume that the MLLM is given an input image $o$ and input text prompt token $r_{0}$, the aim of the reasoning is to obtain the prediction probability $P(y|o,r_{0})$, where $y$ is the final output text token. In general, to utilize CoT, the reasoning problem of predicting $P(y|o,r_{0})$ can be transformed into intermediate steps of predicting $P(z_{t}|o,z_{t-1}),t\in 1,2,\dots,T$. Each $z_{t}=\{r_{t},c_{t}\}$ denotes the augmented textual prompts $r_{t}$ and visual information $c_{t}$ for deepening scene understanding and pushing forward reasoning. After the final reasoning step at $T$, the final output is set to $y=r_{T}$. The target transforms into predicting the following joint probability:
@@ 3.2 Closed-Loop Reasoning Formulation
where $z_{t_{1}:t_{2}}=\{z_{t_{1}},z_{t_{1}+1},\dots,z_{t_{2}}\}$ and $z_{0}=r_{0}$. What is essential in CoT design is to make each step simple enough to match the reasoning capacity of MLLMs.
In imagination space, $\hat{o}$ are images rendered from the space for MLLMs to perceive. We also introduce a set of operators to modify the imagined scenes into new ones. Denote by $v_{t}=\{\hat{o}_{t},r_{0:t}\}$ and $a_{t}$ the operator in step $t$. Our method transforms Eq. 1 into
@@
Eq. 3 shows that one-step reasoning is factorized into a decision function $\pi(a_{t}|v_{t-1})$, a scene modification function $\phi(\hat{o}_{t}|\hat{o}_{t-1},a_{t})$, and a reasoning function $\omega(r_{t}|\hat{o}_{t},r_{1:r-1})$. Given the current imagined scene $\hat{o}_{t-1}$ and the augmented text prompt token $r_{t-1}$, the decision function chooses a specific scene modification operator $a_{t}$. The scene modification function then utilizes $a_{t}$ to update $\hat{o}_{t-1}$ into $\hat{o}_{t}$. The reasoning function finally updates $r_{t-1}$ into $r_{t}$ based on $\hat{o}_{t}$. The decision and reasoning functions are purely based on the native reasoning ability of MLLMs. The scene modification function is implemented inside the imagination space.
Comparing Eq. 1 and Eq. 2, the imagination space transforms CoT reasoning into a closed-loop decision-making and reasoning process. During the reasoning process, the visual state is gradually simplified or approaching a target visual state: the reasoning step at step $t$ is only dependent on $\hat{o}_{t-1},\hat{o}_{t}$ but independent of all previous visual states, leads to effective task decomposition.